# Theoretical Insights into Overparameterized Models in Multi-Task and Replay-Based Continual Learning

### v1

@@ 6 Conclusion
We studied overparameterized linear models in the MTL and CL settings and derived exact theoretical characterizations to better understand the impact of various system parameters over the average generalization error and knowledge transfer. We also analyzed replay-based continual learning and provided theoretical results to describe the generalization error and forgetting of such methods. We also performed extensive experiments with DNNs to show that similar behaviors are observable in the deep MTL and CL models, and understanding the linear overparameterized models can guide our intuition when studying more complex models. Please also consider Appendix H for a discussion on the limitations and future work.
Appendix A Prelude: Single-Task Learning
In this section, we review the theoretical results on the single-task learners as a prelude to our work. The following theorem analyses the error of a single-task learner:
Theorem A.1.
$\mathbb{E}[\mathcal{L}_{t}(w_{t})]=\frac{p\sigma^{2}}{n_{t}-p-1},$ (18)
where the expectation is due to randomness of $X_{t}$ and $z_{t}$.
This theorem resembles the double descent phenomena in neural networks where the error increases by increasing $p$ in the underparameterized regime until the training interpolation threshold $p\approx n_{t}$ is hit. After this point, the training error becomes zero and the model perfectly overfits the training noise. However, the model is benignly overfitted, meaning that increasing $p$ will reduce the noise effect ($\frac{n_{t}\sigma^{2}}{p-n_{t}-1}$ converges to 00 as $p\to\infty$).
Appendix B Naive-Sequential Learner
In this section, we present our theoretical results on the naive-sequential learner of sequentially optimizing Equation 5 on the last train set and ignoring the earlier tasks. We will use this simple theorem as a baseline for comparison with more sophisticated continual learning methods discussed in the coming sections. For simplicity, we assume that $n_{t}=n$ for all tasks.
Theorem B.1.
For a naive sequential learner described in Equation 5, when $p\geq n+2$, it holds that:
$\begin{split}\mathbb{E}[G(w_{T})]&=\frac{n}{Tp}\sum_{i=1}^{T}\|w_{T}^{*}-w_{i}^{*}\|^{2}+\frac{1}{T}(1-\frac{n}{p})\sum_{i=1}^{T}\|w_{i}^{*}\|^{2}+\frac{n\sigma^{2}}{p-n-1},\end{split}$ (20)
$\begin{split}\mathbb{E}[F(w_{T})]=\frac{n}{p(T-1)}\sum_{i=1}^{T-1}\|w_{T}^{*}-w_{i}^{*}\|^{2}.\end{split}$ (21)
As this theorem suggests, even in the most naive sequential learner, increasing the number of parameters $p$ can reduce the forgetting, as also reported by Goldfarb & Hand (2023). However, the generalization error does not behave monotonically with the number of parameters and highly depends on the similarity of the tasks (refer to the $m=0$ curve of the Figures 1(c) and 1(d) for a visualization).
Appendix C Proof of Theorems.
C.1 Useful Lemmas
@@
and we assume that the matrix inverses are defined wherever necessary.
@@
We have already introduced some matrix notations in Section 3.1. In this section, we continue to introduce new helpful notations. Remember that for task $t\in[T]$ with $n_{t}$ samples, we had $X_{t}\in\mathbb{R}^{p\times n_{t}}$, $w^{*}_{t}\in\mathbb{R}^{p}$, $z_{t}\in\mathbb{R}^{n_{t}}$ and $y_{t}=X_{t}^{T}w^{*}_{t}+z_{t}$.
Now consider indices $i,j\in[T]$ such that $i<j$ and denote $n_{ij}=\sum_{t=i}^{j}n_{t}$. We introduce the concatenated data matrix as $X_{ij}\in\mathbb{R}^{p\times n_{ij}}$ which is constructed by concatenating data matrices from $X_{i}$ to $X_{j}$ in the second dimension, i.e. $X_{ij}=\begin{bmatrix}X_{i}&X_{i+1}&...&X_{j}\end{bmatrix}$ With this notation, the matrix representation of all training samples is $X_{1T}\in\mathbb{R}^{p\times\bar{n}}$ where $\bar{n}=\sum_{t=1}^{T}n_{t}$. Additionally, we introduce the matrix $X_{0t}\in\mathbb{R}^{p\times\bar{n}}$ which is built from $X_{1T}$ by replacing all entries with zeros except at the columns corresponding to the task $t$. In other words, $X_{0t}=\begin{bmatrix}0_{p\times r}&X_{t}&0_{p\times q}\end{bmatrix}$, where $0_{p\times r}$ is an all zero matrix with size $p\times r$, $r=\sum_{i=1}^{t-1}n_{i}$ and $q=\sum_{i=t+1}^{T}n_{i}$.
Since our study mainly focuses on overparameterized regimes, the data points $X_{t}\in\mathbb{R}^{p\times n_{t}}$ are random matrices with more rows than columns. Therefore, $(X_{t}^{T}X_{t})^{-1}$ exists almost surely and we define $X_{t}^{\dagger}=X_{t}(X_{t}^{T}X_{t})^{-1}$ and the projection matrix $P_{t}=X_{t}(X_{t}^{T}X_{t})^{-1}X_{t}^{T}$. Additionally, let $P_{ij}=X_{ij}(X_{ij}^{T}X_{ij})^{-1}X_{ij}^{T}$, $X_{ij}^{\dagger}=X_{ij}(X_{ij}^{T}X_{ij})^{-1}$, and $P_{0t}=X_{1T}(X_{1T}^{T}X_{1T})^{-1}X_{0t}^{T}$. With this in mind, we provide some useful lemmas.
Lemma C.3.
Let $X_{12}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be the result of concatenating $X_{1}\in\mathbb{R}^{p\times n_{1}}$ and $X_{2}\in\mathbb{R}^{p\times n_{2}}$. Also let $X_{01}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be result of concatenating $X_{1}$ with a zero matrix. Similarly, build $X_{02}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ by concatenating a zero matrix with $X_{2}$. Assuming $p>n_{1}+n_{2}$ and that all inverses exist, it holds that
@@
where $P_{1}=X_{1}(X_{1}^{T}X_{1})^{-1}X_{1}^{T}$, $X^{\dagger}_{1}=X_{1}(X_{1}^{T}X_{1})^{-1}$ and $I$ is the identity matrix with size $p\times p$.
@@
We start by rewriting $X_{12}^{T}X_{12}$ as a block matrix and then finding its inverse:
$(X_{12}^{T}X_{12})^{-1}=\left(\begin{bmatrix}X^{T}_{1}X_{1}&X^{T}_{1}X_{2}\\ X^{T}_{2}X_{1}&X^{T}_{2}X_{2}\end{bmatrix}\right)^{-1}$
Now using Lemma C.1
$(X_{12}^{T}X_{12})^{-1}=$
$\begin{bmatrix}(X^{T}_{1}X_{1})^{-1}+(X^{T}_{1}X_{1})^{-1}X^{T}_{1}X_{2}H^{-1}X^{T}_{2}X_{1}(X^{T}_{1}X_{1})^{-1}&-(X^{T}_{1}X_{1})^{-1}X^{T}_{1}X_{2}H^{-1}\\ -H^{-1}X^{T}_{2}X_{1}(X^{T}_{1}X_{1})^{-1}&H^{-1}\end{bmatrix},$
where $H=X^{T}_{2}X_{2}-X^{T}_{2}X_{1}(X^{T}_{1}X_{1})^{-1}X^{T}_{1}X_{2}=X^{T}_{2}(I-P_{1})X_{2}$. Now, by noticing that $X_{12}=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}$, we can prove what was desired by doing multiplications:
@@
    $P_{12}$ $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\left(\begin{bmatrix}X^{T}_{1}X_{1}&X^{T}_{1}X_{2}\\ X^{T}_{2}X_{1}&X^{T}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}X^{T}_{1}\\ X^{T}_{2}\end{bmatrix}$
    $=P_{1}+P_{1}X_{2}H^{-1}X^{T}_{2}P_{1}-X_{2}H^{-1}X^{T}_{2}P_{1}-P_{1}X_{2}H^{-1}X^{T}_{2}+X_{2}H^{-1}X^{T}_{2}$
    $=P_{1}+(I-P_{1})X_{2}(X_{2}^{T}(I-P_{1})X_{2})^{-1}X_{2}^{T}(I-P_{1})$
@@
    $P_{01}$ $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\left(\begin{bmatrix}X^{T}_{1}X_{1}&X^{T}_{1}X_{2}\\ X^{T}_{2}X_{1}&X^{T}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}X^{T}_{1}\\ 0\end{bmatrix}$
    $=P_{1}+P_{1}X_{2}H^{-1}X^{T}_{2}P_{1}-X_{2}H^{-1}X^{T}_{2}P_{1}$
    $=P_{1}-(I-P_{1})X_{2}(X_{2}^{T}(I-P_{1})X_{2})^{-1}X^{T}_{2}P_{1}$
@@
    $P_{02}$ $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\left(\begin{bmatrix}X^{T}_{1}X_{1}&X^{T}_{1}X_{2}\\ X^{T}_{2}X_{1}&X^{T}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}0\\ X^{T}_{2}\end{bmatrix}$
    $=-P_{1}X_{2}H^{-1}X^{T}_{2}+X_{2}H^{-1}X^{T}_{2}$
    $=(I-P_{1})X_{2}(X_{2}^{T}(I-P_{1})X_{2})^{-1}X^{T}_{2}$
@@
    $P^{T}_{01}P_{01}$ $=\begin{bmatrix}X_{1}&0\end{bmatrix}\left(\begin{bmatrix}X^{T}_{1}X_{1}&X^{T}_{1}X_{2}\\ X^{T}_{2}X_{1}&X^{T}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}X_{1}^{T}\\ 0\end{bmatrix}$
    $=P_{1}+P_{1}X_{2}H^{-1}X^{T}_{2}P_{1}=P_{1}+P_{1}X_{2}(X_{2}^{T}(I-P_{1})X_{2})^{-1}X^{T}_{2}P_{1}$
@@
    $X^{\dagger}_{12}$
    $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\begin{bmatrix}(X^{T}_{1}X_{1})^{-1}+(X^{T}_{1}X_{1})^{-1}X^{T}_{1}X_{2}H^{-1}X^{T}_{2}X^{\dagger}_{1}&-(X^{T}_{1}X_{1})^{-1}X^{T}_{1}X_{2}H^{-1}\\ -H^{-1}X^{T}_{2}X_{1}(X^{T}_{1}X_{1})^{-1}&H^{-1}\end{bmatrix}$
    $=\begin{bmatrix}X^{\dagger}_{1}-(I-P_{1})X_{2}(X^{T}_{2}(I-P_{1})X_{2})^{-1}X^{T}_{2}X^{\dagger}_{1}&(I-P_{1})X_{2}(X^{T}_{2}(I-P_{1})X_{2})^{-1}\end{bmatrix}$
Lemma C.4.
Let $X_{13}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenating $X_{1}\in\mathbb{R}^{p\times n_{1}}$, $X_{2}\in\mathbb{R}^{p\times n_{2}}$ and $X_{3}\in\mathbb{R}^{p\times n_{3}}$. Also let $X_{01},X_{02},X_{03}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenating $X_{1}$, $X_{2}$ and $X_{3}$ with zero matrices such that $X_{01}=\begin{bmatrix}X_{1}&0&0\end{bmatrix}$ , $X_{02}=\begin{bmatrix}0&X_{2}&0\end{bmatrix}$ and $X_{03}=\begin{bmatrix}0&0&X_{3}\end{bmatrix}$. Define $P_{01}=X_{13}(X^{T}_{13}X_{13})^{-1}X^{T}_{01}$ and $P_{03}=X_{13}(X^{T}_{13}X_{13})^{-1}X^{T}_{03}$ Assuming $p>n_{1}+n_{2}+n_{3}$ and that all inverses exist, it holds that
$P^{T}_{01}P_{03}=-P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}(I-p_{1})(I-\hat{P}_{12})X_{3})^{-1}X_{3}^{T}$
where $P_{1}=X_{1}(X_{1}^{T}X_{1})^{-1}X_{1}^{T}$ and we denote $\hat{P}_{12}=X_{2}(X^{T}_{2}(I-P_{1})X_{2})^{-1}X^{T}_{2}(I-P_{1})$.
@@
$P^{T}_{01}P_{03}$ $=X_{01}(X^{T}_{13}X_{13})^{-1}X^{T}_{13}X_{13}(X^{T}_{13}X_{13})^{-1}X^{T}_{03}$
$=X_{01}(X^{T}_{13}X_{13})^{-1}X^{T}_{03}$
$=X_{01}\left(\begin{bmatrix}X^{T}_{1}X_{1}&X^{T}_{1}X_{2}&X^{T}_{1}X_{3}\\ X^{T}_{2}X_{1}&X^{T}_{2}X_{2}&X^{T}_{2}X_{3}\\ X^{T}_{3}X_{1}&X^{T}_{3}X_{2}&X^{T}_{3}X_{3}\end{bmatrix}\right)^{-1}X^{T}_{03}$
Now we use Lemma C.2 to find the matrix inverse. With the notation introduced in that lemma, we are looking for $-X_{1}E^{-1}UR^{-1}X_{3}^{T}$. Thus we have,
$A$ $=J-HE^{-1}F=X_{2}^{T}(I-P_{1})X_{2}$
$B$ $=K-HE^{-1}G=X_{2}^{T}(I-P_{1})X_{3}$
$C$ $=M-LE^{-1}F=X_{3}^{T}(I-P_{1})X_{2}$
$D$ $=N-LE^{-1}G=X_{3}^{T}(I-P_{1})X_{3}$
$U$ $=G-FA^{-1}B=X_{1}^{T}(I-\hat{P}_{12})X_{3}$
$R$ $=D-CA^{-1}B=X_{3}^{T}(I-P_{1})(I-\hat{P}_{12})X_{3}$
$\Rightarrow P^{T}_{01}P_{03}=-X_{1}E^{-1}UR^{-1}X_{3}^{T}=-P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}(I-P_{1})(I-\hat{P}_{12})X_{3})^{-1}X_{3}^{T}$
Lemma C.5.
@@
$X(X^{T}w^{*}-y)=0\Rightarrow w^{*}=(XX^{T})^{-1}Xy$
noticing that $(XX^{T})^{-1}$ exist almost surely if $X$ is a Gaussian random matrix with $p<n$. ∎
Lemma C.6.
@@
$w^{*}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w-w_{0}\|_{2}^{2}\quad\text{s.t.}\quad X^{T}w=y.$
@@
$w^{*},\lambda^{*}=\operatorname*{arg\,min}_{w,\lambda}\frac{1}{2}\|w-w_{0}\|_{2}^{2}+\lambda^{T}(X^{T}w-y).$
@@
$X^{T}w^{*}=y\Rightarrow-X^{T}X\lambda^{*}+X^{T}w_{0}=y\Rightarrow\lambda^{*}=(X^{T}X)^{-1}X^{T}w_{0}-(X^{T}X)^{-1}y$
$\Rightarrow w^{*}=-X(X^{T}X)^{-1}X^{T}w_{0}+X(X^{T}X)^{-1}y+w_{0}=(I-P)w_{0}+X^{\dagger}y$
Lemma C.7.
Assume matrices $X_{1}\in\mathbb{R}^{p\times n_{1}}$ and $X_{2}\in\mathbb{R}^{p\times n_{2}}$ to be random matrices with entries being independently sampled from the standard Gaussian distribution $\mathcal{N}(0,1)$. Assume $p>n_{1}+n_{2}+1$ and let $P_{1}=X_{1}(X^{T}_{1}X_{1})^{-1}X^{T}_{1}$ be the orthogonal projection matrix that projects onto the column space of $X_{1}$. Also assume $X_{12}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be the result of concatenation of $X_{1}$ and $X_{2}$, and $X_{01}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be the result of concatenation of $X_{1}$ with a zero matrix. Additionally, let $P_{12}=X_{12}(X^{T}_{12}X_{12})^{-1}X^{T}_{12}$ and $P_{01}=X_{12}(X^{T}_{12}X_{12})^{-1}X^{T}_{01}$. Assuming $w\in\mathbb{R}^{p}$ is a fixed given vector, the following equalities hold:
@@
    $\mathbb{E}[\|P_{01}w\|^{2}]=\frac{n_{1}}{p}(1+\frac{n_{2}}{p-(n_{1}+n_{2})-1})\|w\|^{2}$

### v2

@@ 6 Conclusion
We studied overparameterized linear models in the MTL and CL settings and derived exact theoretical characterizations to better understand the impact of various system parameters over the average generalization error and knowledge transfer. We also analyzed replay-based continual learning and provided theoretical results to describe the generalization error and forgetting of such methods. We also performed extensive experiments with DNNs to show that similar behaviors are observable in the deep MTL and CL models, and understanding the linear overparameterized models can guide our intuition when studying more complex models. Please see Appendix E for a discussion on the limitations and future work.
Appendix A Proof of Theorems.
A.1 Useful Lemmas
@@
and we assume that the matrix inverses exist wherever necessary.
@@
We have already introduced some matrix notations in Section 3.1. In this section, we continue to introduce new helpful notations. Remember that for task $t\in[T]$ with $n_{t}$ samples, we had $X_{t}\in\mathbb{R}^{p\times n_{t}}$, $w^{*}_{t}\in\mathbb{R}^{p}$, $z_{t}\in\mathbb{R}^{n_{t}}$ and $y_{t}=X_{t}^{\top}w^{*}_{t}+z_{t}$.
Now consider indices $i,j\in[T]$ such that $i<j$ and denote $n_{ij}=\sum_{t=i}^{j}n_{t}$. We introduce the concatenated data matrix as $X_{ij}\in\mathbb{R}^{p\times n_{ij}}$ which is constructed by concatenating data matrices from $X_{i}$ to $X_{j}$ in the second dimension, i.e. $X_{ij}=\begin{bmatrix}X_{i}&X_{i+1}&...&X_{j}\end{bmatrix}$ With this notation, the matrix representation of all training samples is $X_{1:T}\in\mathbb{R}^{p\times\bar{n}}$ where $\bar{n}=\sum_{t=1}^{T}n_{t}$. Additionally, we introduce the matrix $X_{0|t}\in\mathbb{R}^{p\times\bar{n}}$ which is built from $X_{1:T}$ by replacing all entries with zeros except at the columns corresponding to the task $t$. In other words, $X_{0|t}=\begin{bmatrix}0_{p\times r}&X_{t}&0_{p\times q}\end{bmatrix}$, where $0_{p\times r}$ is an all zero matrix with size $p\times r$, $r=\sum_{i=1}^{t-1}n_{i}$ and $q=\sum_{i=t+1}^{T}n_{i}$.
Since our study mainly focuses on overparameterized regimes, the data points $X_{t}\in\mathbb{R}^{p\times n_{t}}$ are random matrices with more rows than columns. Therefore, $(X_{t}^{\top}X_{t})^{-1}$ exists almost surely and we define $X_{t}^{\dagger}=X_{t}(X_{t}^{\top}X_{t})^{-1}$ and the projection matrix $P_{t}=X_{t}(X_{t}^{\top}X_{t})^{-1}X_{t}^{\top}$. Additionally, let $P_{ij}=X_{ij}(X_{ij}^{\top}X_{ij})^{-1}X_{ij}^{\top}$, $X_{ij}^{\dagger}=X_{ij}(X_{ij}^{\top}X_{ij})^{-1}$, and $P_{0|t}=X_{1:T}(X_{1:T}^{\top}X_{1:T})^{-1}X_{0|t}^{\top}$. With this in mind, we provide some useful lemmas. [^2] Notice that $P_{0|t}$, by definition is not a projection matrix and we use this notation just for consistency.
Lemma A.3.
Let $X_{1:2}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be the result of concatenating $X_{1}\in\mathbb{R}^{p\times n_{1}}$ and $X_{2}\in\mathbb{R}^{p\times n_{2}}$. Also let $X_{0|1}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be result of concatenating $X_{1}$ with a zero matrix. Similarly, build $X_{0|2}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ by concatenating a zero matrix with $X_{2}$. Assuming $p>n_{1}+n_{2}$ and that all inverses exist, it holds that
@@
where $P_{1}=X_{1}(X_{1}^{\top}X_{1})^{-1}X_{1}^{\top}$, $X^{\dagger}_{1}=X_{1}(X_{1}^{\top}X_{1})^{-1}$ and $I$ is the identity matrix with size $p\times p$.
@@
We start by rewriting $X_{1:2}^{\top}X_{1:2}$ as a block matrix and then finding its inverse:
$(X_{1:2}^{\top}X_{1:2})^{-1}=\left(\begin{bmatrix}X^{\top}_{1}X_{1}&X^{\top}_{1}X_{2}\\ X^{\top}_{2}X_{1}&X^{\top}_{2}X_{2}\end{bmatrix}\right)^{-1}$
Now using Lemma A.1
$(X_{1:2}^{\top}X_{1:2})^{-1}$ $=\begin{bmatrix}(X^{\top}_{1}X_{1})^{-1}+(X^{\top}_{1}X_{1})^{-1}X^{\top}_{1}X_{2}H^{-1}X^{\top}_{2}X_{1}(X^{\top}_{1}X_{1})^{-1}&-(X^{\top}_{1}X_{1})^{-1}X^{\top}_{1}X_{2}H^{-1}\\ -H^{-1}X^{\top}_{2}X_{1}(X^{\top}_{1}X_{1})^{-1}&H^{-1}\end{bmatrix},$
where $H=X^{\top}_{2}X_{2}-X^{\top}_{2}X_{1}(X^{\top}_{1}X_{1})^{-1}X^{\top}_{1}X_{2}=X^{\top}_{2}(I-P_{1})X_{2}$. Now, by noticing that $X_{1:2}=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}$, we can prove what was desired by doing multiplications:
@@
    $P_{1:2}$ $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\left(\begin{bmatrix}X^{\top}_{1}X_{1}&X^{\top}_{1}X_{2}\\ X^{\top}_{2}X_{1}&X^{\top}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}X^{\top}_{1}\\ X^{\top}_{2}\end{bmatrix}$
    $=P_{1}+P_{1}X_{2}H^{-1}X^{\top}_{2}P_{1}-X_{2}H^{-1}X^{\top}_{2}P_{1}-P_{1}X_{2}H^{-1}X^{\top}_{2}+X_{2}H^{-1}X^{\top}_{2}$
    $=P_{1}+(I-P_{1})X_{2}(X_{2}^{\top}(I-P_{1})X_{2})^{-1}X_{2}^{\top}(I-P_{1})$
@@
    $P_{0|1}$ $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\left(\begin{bmatrix}X^{\top}_{1}X_{1}&X^{\top}_{1}X_{2}\\ X^{\top}_{2}X_{1}&X^{\top}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}X^{\top}_{1}\\ 0\end{bmatrix}$
    $=P_{1}+P_{1}X_{2}H^{-1}X^{\top}_{2}P_{1}-X_{2}H^{-1}X^{\top}_{2}P_{1}$
    $=P_{1}-(I-P_{1})X_{2}(X_{2}^{\top}(I-P_{1})X_{2})^{-1}X^{\top}_{2}P_{1}$
@@
    $P_{0|2}$ $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\left(\begin{bmatrix}X^{\top}_{1}X_{1}&X^{\top}_{1}X_{2}\\ X^{\top}_{2}X_{1}&X^{\top}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}0\\ X^{\top}_{2}\end{bmatrix}$
    $=-P_{1}X_{2}H^{-1}X^{\top}_{2}+X_{2}H^{-1}X^{\top}_{2}$
    $=(I-P_{1})X_{2}(X_{2}^{\top}(I-P_{1})X_{2})^{-1}X^{\top}_{2}$
@@
    $P^{\top}_{0|1}P_{0|1}$ $=\begin{bmatrix}X_{1}&0\end{bmatrix}\left(\begin{bmatrix}X^{\top}_{1}X_{1}&X^{\top}_{1}X_{2}\\ X^{\top}_{2}X_{1}&X^{\top}_{2}X_{2}\end{bmatrix}\right)^{-1}\begin{bmatrix}X_{1}^{\top}\\ 0\end{bmatrix}$
    $=P_{1}+P_{1}X_{2}H^{-1}X^{\top}_{2}P_{1}=P_{1}+P_{1}X_{2}(X_{2}^{\top}(I-P_{1})X_{2})^{-1}X^{\top}_{2}P_{1}$
@@
    $X^{\dagger}_{1:2}$ $=\begin{bmatrix}X_{1}&X_{2}\end{bmatrix}\begin{bmatrix}(X^{\top}_{1}X_{1})^{-1}+(X^{\top}_{1}X_{1})^{-1}X^{\top}_{1}X_{2}H^{-1}X^{\top}_{2}X^{\dagger}_{1}&-(X^{\top}_{1}X_{1})^{-1}X^{\top}_{1}X_{2}H^{-1}\\ -H^{-1}X^{\top}_{2}X_{1}(X^{\top}_{1}X_{1})^{-1}&H^{-1}\end{bmatrix}$
    $=\begin{bmatrix}X^{\dagger}_{1}-(I-P_{1})X_{2}(X^{\top}_{2}(I-P_{1})X_{2})^{-1}X^{\top}_{2}X^{\dagger}_{1}&(I-P_{1})X_{2}(X^{\top}_{2}(I-P_{1})X_{2})^{-1}\end{bmatrix}$
Lemma A.4.
Let $X_{1:3}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenating $X_{1}\in\mathbb{R}^{p\times n_{1}}$, $X_{2}\in\mathbb{R}^{p\times n_{2}}$ and $X_{3}\in\mathbb{R}^{p\times n_{3}}$. Also let $X_{0|1},X_{0|2},X_{0|3}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenating $X_{1}$, $X_{2}$ and $X_{3}$ with zero matrices such that $X_{0|1}=\begin{bmatrix}X_{1}&0&0\end{bmatrix}$ , $X_{0|2}=\begin{bmatrix}0&X_{2}&0\end{bmatrix}$ and $X_{0|3}=\begin{bmatrix}0&0&X_{3}\end{bmatrix}$. Define $P_{0|1}=X_{1:3}(X^{\top}_{1:3}X_{1:3})^{-1}X^{\top}_{0|1}$ and $P_{0|3}=X_{1:3}(X^{\top}_{1:3}X_{1:3})^{-1}X^{\top}_{0|3}$ Assuming $p>n_{1}+n_{2}+n_{3}$ and that all inverses exist, it holds that
$P^{\top}_{0|1}P_{0|3}=-P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}(I-P_{1})(I-\hat{P}_{1:2})X_{3})^{-1}X_{3}^{\top}$
where $P_{1}=X_{1}(X_{1}^{\top}X_{1})^{-1}X_{1}^{\top}$ and we denote $\hat{P}_{1:2}=X_{2}(X^{\top}_{2}(I-P_{1})X_{2})^{-1}X^{\top}_{2}(I-P_{1})$.
@@
$P^{\top}_{0|1}P_{0|3}$ $=X_{0|1}(X^{\top}_{1:3}X_{1:3})^{-1}X^{\top}_{1:3}X_{1:3}(X^{\top}_{1:3}X_{1:3})^{-1}X^{\top}_{0|3}$
$=X_{0|1}(X^{\top}_{1:3}X_{1:3})^{-1}X^{\top}_{0|3}$
$=X_{0|1}\left(\begin{bmatrix}X^{\top}_{1}X_{1}&X^{\top}_{1}X_{2}&X^{\top}_{1}X_{3}\\ X^{\top}_{2}X_{1}&X^{\top}_{2}X_{2}&X^{\top}_{2}X_{3}\\ X^{\top}_{3}X_{1}&X^{\top}_{3}X_{2}&X^{\top}_{3}X_{3}\end{bmatrix}\right)^{-1}X^{\top}_{0|3}$
Now we use Lemma A.2 to find the matrix inverse. With the notation introduced in that lemma, we are looking for $-X_{1}E^{-1}UR^{-1}X_{3}^{\top}$. Thus we have,
$A$ $=J-HE^{-1}F=X_{2}^{\top}(I-P_{1})X_{2}$
$B$ $=K-HE^{-1}G=X_{2}^{\top}(I-P_{1})X_{3}$
$C$ $=M-LE^{-1}F=X_{3}^{\top}(I-P_{1})X_{2}$
$D$ $=N-LE^{-1}G=X_{3}^{\top}(I-P_{1})X_{3}$
$U$ $=G-FA^{-1}B=X_{1}^{\top}(I-\hat{P}_{1:2})X_{3}$
$R$ $=D-CA^{-1}B=X_{3}^{\top}(I-P_{1})(I-\hat{P}_{1:2})X_{3}$
$\Rightarrow P^{\top}_{0|1}P_{0|3}=-X_{1}E^{-1}UR^{-1}X_{3}^{\top}=-P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}(I-P_{1})(I-\hat{P}_{1:2})X_{3})^{-1}X_{3}^{\top}$
Lemma A.5.
@@
$X(X^{\top}w^{*}-y)=0\Rightarrow w^{*}=(XX^{\top})^{-1}Xy$
noticing that $(XX^{\top})^{-1}$ exist almost surely if $X$ is a Gaussian random matrix with $p<n$. ∎
Lemma A.6.
@@
$w^{*}=\operatorname*{arg\,min}_{w\in\mathbb{R}^{p}}\|w-w_{0}\|_{2}^{2}\quad\text{s.t.}\quad X^{\top}w=y.$
@@
$w^{*},\lambda^{*}=\operatorname*{arg\,min}_{w,\lambda}\frac{1}{2}\|w-w_{0}\|_{2}^{2}+\lambda^{\top}(X^{\top}w-y).$
@@
$X^{\top}w^{*}=y\Rightarrow-X^{\top}X\lambda^{*}+X^{\top}w_{0}=y\Rightarrow\lambda^{*}=(X^{\top}X)^{-1}X^{\top}w_{0}-(X^{\top}X)^{-1}y$
$\Rightarrow w^{*}=-X(X^{\top}X)^{-1}X^{\top}w_{0}+X(X^{\top}X)^{-1}y+w_{0}=(I-P)w_{0}+X^{\dagger}y$
Lemma A.7.
Assume matrices $X_{1}\in\mathbb{R}^{p\times n_{1}}$ and $X_{2}\in\mathbb{R}^{p\times n_{2}}$ to be random matrices with entries being independently sampled from the standard Gaussian distribution $\mathcal{N}(0,1)$. Assume $p>n_{1}+n_{2}+1$ and let $P_{1}=X_{1}(X^{\top}_{1}X_{1})^{-1}X^{\top}_{1}$ be the orthogonal projection matrix that projects onto the column space of $X_{1}$. Also assume $X_{1:2}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be the result of concatenation of $X_{1}$ and $X_{2}$, and $X_{0|1}\in\mathbb{R}^{p\times(n_{1}+n_{2})}$ be the result of concatenation of $X_{1}$ with a zero matrix. Additionally, let $P_{1:2}=X_{1:2}(X^{\top}_{1:2}X_{1:2})^{-1}X^{\top}_{1:2}$ and $P_{0|1}=X_{1:2}(X^{\top}_{1:2}X_{1:2})^{-1}X^{\top}_{0|1}$. Assuming $w\in\mathbb{R}^{p}$ is a fixed given vector, the following equalities hold:
@@
    $\mathbb{E}[\|P_{0|1}w\|^{2}]=\frac{n_{1}}{p}(1+\frac{n_{2}}{p-(n_{1}+n_{2})-1})\|w\|^{2}$